Shift report 20010826 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Aug 26 2001 - 00:01:03 EDT

  • Next message: Ian Bearden: "RE: very preliminary brahms dndeta with model prediction"

    ------------------------------------------------------------
    Date: Sun Aug 26 00:01:02 2001, Shift: 16:00-24:00
    Supervisor: Alexandru JIPA, with: Asbjoern HANSSEN
    Log book#8 pages 23 - 24
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beams up to 21.40. Beams dumping up to 22.26. Fast decrease in Blue Beam intensity
    2. Three good runs (4402, 4407, 4408; run 4408 still works); more than 400K events
    3. At T5 at the end of the run 4407 the anodes colour was blue.
    Some problems with IoServer between run 4402 and run 4407. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Aug 26 2001 - 00:01:38 EDT